# How hydrodynamic interactions alter polymer stretching in turbulence

### v1

@@ Abstract
Hydrodynamic interactions (HI) between segments of a polymer have long been known to strongly affect polymer stretching in laminar viscometric flows. Yet the role of HI in fluctuating turbulent flows remains unclear. Using Brownian dynamics simulations, we examine the stretching dynamics of bead-spring chains with inter-bead HI, as they are transported in a homogeneous isotropic turbulent flow (within the ultra-dilute, one-way coupling regime). We find that HI-endowed chains exhibit a steeper coil-stretch transition as the elastic relaxation time is increased, i.e., HI cause less stretching of stiff polymers and more stretching of moderately to highly elastic polymers. The probability distribution function of the end-to-end extension is also modified, with HI significantly limiting the range of extensions over which a power-law range appears. On quantifying the repeated stretching and recoiling of chains by computing persistence time distributions, we find that HI delays migration between stretched and coiled states. These effects of HI, which are consistent with chains experiencing an effective conformation-dependent drag, are sensitive to the level of coarse-graining in the bead-spring model. Specifically, an HI-endowed dumbbell, which cannot form a physical coil, is unable to experience the hydrodynamic shielding effect of HI. Our results highlight the importance of incorporating an extension-dependent drag force in dumbbell-based simulations of turbulent polymer solutions. To develop and test such an augmented dumbbell model, we propose the use of a time-correlated Gaussian random flow, in which the turbulent stretching statistics are shown to be well-approximated.
@@ I Introduction
A dissolved polymer when stretched exerts elastic feedback forces onto the solvent. At the macroscale, this feedback renders the solution viscoelastic and produces phenomena like turbulent drag reduction [1, 2, 3, 4] and inertialess chaotic flow [5, 6]. At the microscale, that is, at the scale of the polymer, the disturbance flow produced by the feedback force of one part of the polymer can alter the motion of all other parts. Such solvent-mediated interactions between different parts of the polymer are termed hydrodynamic interactions (HI), and their influence on the stretching of polymers has been studied extensively in laminar viscometric flows [7, 8, 9, 10, 11, 12, 13]. In extension-dominated flows, HI are known to give rise to an effective conformation-dependent drag [14, 15, 16, 17, 18, 19, 8]: When coiled, the inner portions of the polymer experience a much smaller drag than would be expected based on the undisturbed flow; this shielding effect weakens as the polymer stretches out; the result is that coiled (and stretched) polymers are more likely to remain coiled (and stretched) in the presence of HI [8, 12]. In steady homogeneous extensional flows, this effect gives rise to a coil-stretch hysteresis [14, 15, 19, 8, 20] as the Weissenberg number ($\mathrm{Wi}$)—the product of the strain-rate and the polymer’s largest elastic relaxation time—is varied. The consequences of HI for polymer stretching in turbulence, however, are not well understood. With rare exceptions [21, 22, 23, 24], HI has typically been ignored in studies of the Lagrangian dynamics of polymers in turbulent flows.
@@
In the context of turbulence one is usually interested in situations where polymers are highly stretched, so that viscoelastic effects are strong enough to produce drag reduction [2, 3, 4] in high Reynolds number flows, or generate elasto-inertial [25, 26, 27, 28] and elastic turbulence [29, 6, 30, 31, 32, 33] in moderate and low Reynolds number flows. Will HI be relevant in such scenarios, given that the effects of HI are relatively weak for fully-extended polymers? Yes, because in turbulence the distribution of polymer extensions is always broad, owing to the fluctuating nature of the strain-rate.
@@
Indeed, the theory of Balkovsky, Fouxon, and Lebedev [34, 35] and Chertkov [36], developed for a dumbbell (without HI) in a general chaotic flow, predicts that the probability distribution function (PDF) of the extension $R$ will always exhibit a power-law behaviour. The corresponding exponent increases with $\mathrm{Wi}$, crossing 1-1 at $\mathrm{Wi}=1/2$, when $\mathrm{Wi}$ is defined as the product of the polymer relaxation time and the Lyapunov exponent of Lagrangian trajectories in the flow. Since a dependence of $R^{-1}$ implies a non-normalizable PDF, in the idealized case of a Hookean dumbbell, $\mathrm{Wi}=1/2$ is treated as the critical point of a coil-stretch transition. This is not mere convention—the PDF of $R$ does indeed exhibit hallmark features of phase transitions near $\mathrm{Wi}=1/2$, such as a slowing down of dynamics [37, 38, 39] and a maximization of entropy [40, 41]. The broad, power-law PDF of $R$, which has been measured in experiments [42, 43, 44] and clearly observed in simulations of free-draining (no HI) dumbbells and chains [45, 46, 47, 4, 39], implies that many polymers will be coiled even at large $\mathrm{Wi}$. When a stretched polymer is advected into a region of the flow with low straining, it will recoil and remain coiled until it encounters high strain-rates (the residence time in coiled states follows a Poisson distribution [39]). This is when HI will be important. Indeed, we show here that HI delays the migration between coiled and stretched states, and thereby alters qualitatively the PDF of extension.
@@
In this work, we study the effects of HI on polymer stretching in homogeneous isotropic turbulence, using a bead-spring chain model for the polymer. For simplicity, we ignore the possibility of excluded volume (EV) interactions between the beads and focus on HI alone. We first describe the bead-spring model in Sec. II, and discuss how models with different levels of coarse-graining—from dumbbells to many-bead chains—are compared. The turbulent carrier flow is also discussed in this section. We then show that the effects of HI are qualitatively different for dumbbells and chains (Sec. III). Since a dumbbell cannot bend and form a coil, it does not experience the shielding effect of HI and, in contrast to a chain, is stretched out more easily in the presence of HI. (The effects of HI on dumbbells in a fluctuating flow, including the enhancement of stretching, have been elucidated in recent analytical work by Picardo and Vincenzi [24]). In Sec. IV, we show that HI steepens the coil stretch transition of chains, in a manner than cannot be captured by a simple rescaling of $\mathrm{Wi}$. Moreover, the PDF of $R$ is found to be modified at small and intermediate values of $R$, such that the power-law range is significantly reduced. By analyzing the persistence time of chains in stretched and coiled configurations, we directly confirm that these effects of HI are a consequence of HI-endowed chains taking longer to migrate between states of small and large extensions. Finally, we check whether the effects of HI on turbulent stretching can be predicted using a random model for the turbulent velocity gradient. Our results, in Sec. V, show that a time-correlated Gaussian random model works well, suggesting its use as a testing ground for developing coarse-grained polymer models, capable of capturing the effects of HI in fluctuating flows. We end in Sec. VI with a summary of our results and suggestions for future work.
@@ II.1 Bead-spring chain with hydrodynamic interactions
${\rm d}{\bm{X}}_{c}$ $=\bm{u}(\bm{X}_{c}(t),t)\;{\rm d}t+\dfrac{1}{N_{\mathrm{b}}}\sqrt{\frac{Q_{\mathrm{eq}}^{2}}{6\tau_{s}}}\sum_{i=1}^{N_{\mathrm{b}}}{\rm d}\bm{W}_{i}(t),$ (1)
${\rm d}{\bm{Q}_{i}}$ $=\bm{\kappa}(t)\cdot\bm{Q}_{i}\;{\rm d}t+\dfrac{1}{4\tau_{s}}\sum_{j=1}^{N_{\mathrm{b}}}\left(\bm{D}_{i+1,j}-\bm{D}_{i,j}\right)\cdot\bm{F}_{j}^{\rm E}\;{\rm d}t+\sqrt{\dfrac{Q_{\mathrm{eq}}^{2}}{6\tau_{s}}}\sum_{j=1}^{i+1}\left(\bm{B}_{i+1,j}-\bm{B}_{i,j}\right)\cdot{\rm d}\bm{W}_{j}(t).$ (2)
@@
The links between neighboring beads are phantom FENE springs with spring constant $H$. If $\zeta$ is the Stokes drag coefficient of the beads, then each link is characterized by an elastic time scale $\tau_{s}=\zeta/4H$. The net spring force exerted on bead $i$ is $\bm{F}_{i}^{\rm E}=f_{i}\bm{Q}_{i}-f_{i-1}\bm{Q}_{i-1}$, where the FENE interaction coefficients $f_{i}=\left({1-|\bm{Q}_{i}|^{2}/Q_{\mathrm{m}}^{2}}\right)^{-1}$ ensure that the extension of each link does not exceed its maximum length $Q_{\mathrm{m}}$. The contour length or maximum length of the chain is then given by $R_{\mathrm{m}}=(N_{\mathrm{b}}-1)Q_{\mathrm{m}}$. [Note that in Eq. (2), one must set $\bm{Q}_{0}=\bm{Q}_{N_{\mathrm{b}}}=0$ in the equations for ${\rm d}\bm{Q}_{1}$} and ${\rm d}{\bm{Q}}_{N_{\mathrm{b}}-1}$.]
The Brownian forces that act on the beads are represented by independent increments of vectorial Wiener processes ${\rm d}\bm{W}_{i}(t)$. The competition between Brownian and elastic forces sets the equilibrium length scale $Q_{\mathrm{eq}}=\sqrt{3k_{B}T/H}$, where $k_{B}$ is the Boltzmann constant and $T$ is the temperature. $Q_{\mathrm{eq}}$ is the root-mean-square (r.m.s.) extension of the link at equilibrium in a still fluid, in the absence of inter-bead interactions, and in the Hookean limit of $Q_{\mathrm{eq}}\ll Q_{\mathrm{m}}$. The corresponding r.m.s. end-to-end extension of the chain at equilibrium is $R_{\mathrm{eq}}=Q_{\mathrm{eq}}\sqrt{N_{\mathrm{b}}-1}$ [48].
@@
Equations (1) to (7) constitute the bead-spring chain model with hydrodynamic interactions. For a given number of beads, $N_{\mathrm{b}}$, the chain is parameterized by $Q_{\mathrm{eq}}$, $Q_{\mathrm{m}}$, $\tau_{s}$, and the HI parameter $h^{\star}$ (setting $h^{\star}=0$ yields the free-draining Rouse model [48, 49]).
II.2 Brownian dynamics simulations

### v2

@@ Abstract
Hydrodynamic interactions (HI) between segments of a polymer have long been known to strongly affect polymer stretching in laminar viscometric flows. Yet the role of HI in fluctuating turbulent flows remains unclear. Using Brownian dynamics simulations, we examine the stretching dynamics of bead-spring chains with inter-bead HI, as they are transported in a homogeneous isotropic turbulent flow (within the ultra-dilute, one-way coupling regime). We find that the effects of HI are negligibly small for dumbbells but become increasingly prominent as the polymer model is refined, i.e., as the number of beads are increased. HI-endowed chains exhibit a steeper coil-stretch transition as the elastic relaxation time is increased, i.e., HI cause less stretching of stiff polymers and more stretching of moderately elastic polymers. The probability distribution function of the end-to-end extension is also modified, with HI significantly limiting the range of extensions over which a power-law range appears. On quantifying the repeated stretching and recoiling of chains by computing persistence time distributions, we find that HI delays migration between stretched and coiled states. These effects of HI, which are consistent with chains experiencing an effective conformation-dependent drag, are sensitive to the level of coarse-graining in the bead-spring model. Specifically, an HI-endowed dumbbell, which cannot form a physical coil, is unable to experience the hydrodynamic shielding effect of HI. Our results highlight the importance of incorporating an extension-dependent drag force in dumbbell-based simulations of turbulent polymer solutions. To develop and test such an augmented dumbbell model, we propose the use of a time-correlated Gaussian random flow, in which the turbulent stretching statistics are shown to be well-approximated.
@@ I Introduction
A dissolved polymer when stretched exerts elastic feedback forces onto the solvent. At the macroscale, this feedback renders the solution viscoelastic and produces phenomena like turbulent drag reduction [1, 2, 3, 4] and inertialess chaotic flow [5, 6]. At the microscale, that is, at the scale of the polymer, the disturbance flow produced by the feedback force of one part of the polymer can alter the motion of all other parts. Such solvent-mediated interactions between different parts of the polymer are termed hydrodynamic interactions (HI). Their consequences for polymer stretching and stress have been studied extensively in laminar viscometric and extension-dominated flows [7, 8, 9, 10, 11, 12, 13]. The effects of HI in turbulent flows, however, are not well understood. With rare exceptions [14, 15, 16, 17], HI have typically been ignored in studies of the Lagrangian dynamics of polymers in turbulence. Our goal in this work is to examine the nature and extent of the influence of HI on the stretching of polymers in turbulent fluctuating flows, while considering a hierarchy of bead-spring polymer models.
To set the stage, it is helpful to recall the physical consequences of HI. A long flexible polymer may be pictured as a freely-jointed chain of $N_{\textsc{k}}$ Kuhn segments, each of length $b_{\textsc{k}}$. Under quiescent (thermodynamic) equilibrium conditions, the chain forms a random coil with size $R_{\rm eq}$. When such a coil is dragged through a solvent, interior segments are partially shielded from the background flow because the disturbance flows created by many segments superpose. In effect, the solvent cannot easily “drain” through the coil at the same rate that it would without HI. Hence, the drag experienced by the coil is akin to that of a sphere with an effective hydrodynamic size. As the polymer stretches out, the geometric configuration changes, shielding weakens, and the effective drag changes accordingly. In extension-dominated laminar flows, this produces an effective conformation-dependent drag [18, 19, 20, 21, 22, 23, 8], which in turn can yield coil-stretch hysteresis [18, 19, 23, 8, 24]: coiled configurations tend to remain coiled longer, and stretched configurations tend to remain stretched longer [8, 12]. These are not subtle effects in strong extension; they can qualitatively reorganize stretching dynamics by slowing migration between conformational states, and lead to hysteresis in rheological measurements [25, 26, 27].
@@
In the context of turbulence one is usually interested in situations where polymers are highly stretched, so that viscoelastic effects are strong enough to produce drag reduction [1, 2, 3, 4] in high Reynolds number flows, or generate elasto-inertial [28, 29, 30, 31] and elastic turbulence [32, 6, 33, 34, 35, 36] in moderate and low Reynolds number flows. Will HI be relevant in such scenarios, given that the effects of HI are relatively weak for fully-extended polymers? Yes, because in turbulence the distribution of polymer extensions is always broad, owing to the fluctuating nature of the strain-rate.
@@
Indeed, the theory of Balkovsky, Fouxon, and Lebedev [37, 38] and Chertkov [39], developed for a dumbbell (without HI) in a general chaotic flow, predicts that the probability distribution function (PDF) of the extension $R$ will always exhibit a power-law behaviour. The corresponding exponent increases with $\mathrm{Wi}$, crossing 1-1 at $\mathrm{Wi}=1/2$, when $\mathrm{Wi}$ is defined as the product of the polymer relaxation time and the Lyapunov exponent of Lagrangian trajectories in the flow. Since a dependence of $R^{-1}$ implies a non-normalizable PDF, in the idealized case of a Hookean dumbbell, $\mathrm{Wi}=1/2$ is treated as the critical point of a coil-stretch transition. This is not mere convention—the PDF of $R$ does indeed exhibit hallmark features of phase transitions near $\mathrm{Wi}=1/2$, such as a slowing down of dynamics [40, 41, 42] and a maximization of entropy [43, 44]. The broad, power-law PDF of $R$, which has been measured in experiments [45, 46, 47] and clearly observed in simulations of free-draining (no HI) dumbbells and chains [48, 49, 50, 4, 42], implies that many polymers will be coiled even at large $\mathrm{Wi}$. When a stretched polymer is advected into a region of the flow with low straining, it will recoil and remain coiled until it encounters high strain-rates (the residence time in coiled states follows a Poisson distribution [42]). This is when HI will be important. Indeed, we show here that HI delay the migration between coiled and stretched states, and thereby alter the PDF of extension qualitatively.
In this work, we study the effects of HI on polymer stretching in homogeneous isotropic turbulence, using a bead-spring chain model for the polymer. Ideally, one would like to simulate the polymer at Kuhn-segment resolution using a bead-rod chain; however, such simulations with HI in turbulence are impractical at present, especially when large ensembles of long trajectories are required. We therefore resort to the coarse-grained bead-spring description in which the bead-rod chain is divided into $N_{\mathrm{b}}$ subchains, each represented by an isotropic “bead” connected by entropic springs.
In the Rouse model, each bead has an isotropic Stokes friction coefficient $\zeta$, and the HI between different beads are neglected. As a result, the overall friction of a Rouse chain is simply $N_{\mathrm{b}}\,\zeta$ and the drag on the chain is independent of the chain’s conformation. When one retains inter-bead HI, for example via the Rotne–Prager–Yamakawa (RPY) mobility tensor, then the model acquires the ability to represent shielding at the scale of bead-bead separations. The chain will therefore experience a conformation-dependent drag, owing to shielding at equilibrium and deshielding as the chain uncoils and stretches out. Importantly, the extent to which a coarse-grained model can represent this conformation dependence varies with $N_{\mathrm{b}}$: a few widely spaced beads may under-resolve shielding effects that a finer representation (many closely spaced beads) can capture. This connection between the effects of HI and coarse-graining is well established for simple rheometric flows [51, 52, 11], wherein it is clear that a dumbbell with HI fails to capture the effects of HI on a more realistic, multi-bead chain. Therefore, to determine the influence of HI on polymer stretching in turbulence, it is not sufficient to merely add HI to a dumbbell model. A weak effect of HI on dumbbell stretching (a result obtained in recent analytical work [17] and supported by our simulations) need not imply that HI is unimportant but only that a more refined polymer model (with more beads) may be needed to reveal the full effects of HI. Our work shows that the latter is true.
@@
The bead-spring chain used in our study is described in Sec. II. For simplicity, we ignore excluded-volume interactions and focus on HI alone. In this section, we describe how the parameters of the chain are varied with $N_{\mathrm{b}}$ in order for chains with increasing $N_{\mathrm{b}}$ to correspond to increasingly refined descriptions of the same physical polymer. We also discuss the turbulent carrier flow here. In Sec. III, we show that HI effects are qualitatively different for dumbbells and multibead chains: a dumbbell cannot bend into a physical coil and therefore does not experience shielding in the same manner as a chain; accordingly, weak sensitivity to HI at $N_{\mathrm{b}}=2$ does not imply a weak sensitivity at higher $N_{\mathrm{b}}$. In Sec. IV we show that HI steepens the coil-stretch transition for chains and qualitatively modifies the extension PDF at small and intermediate $R$, reducing the range over which a clear power-law is visible. By analyzing the persistence time of chains in stretched and coiled configurations, we directly confirm that these effects of HI are a consequence of HI-endowed chains taking longer to migrate between states of small and large extensions. Finally, we check whether the effects of HI on turbulent stretching can be predicted using a random model for the turbulent velocity gradient. Our results, in Sec. V, show that a time-correlated Gaussian random model works well, suggesting its use as a testing ground for developing coarse-grained polymer models, capable of capturing the effects of HI in fluctuating flows. We end in Sec. VI where we discuss the implications of our results for the simulation of viscoelastic turbulence and suggest directions for future work.
@@ II.1 Bead-spring chain with hydrodynamic interactions
${\rm d}{\bm{X}}_{c}$ $=\bm{u}(\bm{X}_{c}(t),t)\;{\rm d}t+\dfrac{1}{N_{\mathrm{b}}}\sqrt{\frac{Q_{\mathrm{eq}}^{2}}{6\tau_{\mathrm{s}}}}\sum_{i=1}^{N_{\mathrm{b}}}{\rm d}\bm{W}_{i}(t),$ (1)
${\rm d}{\bm{Q}_{i}}$ $=\bm{\kappa}(t)\cdot\bm{Q}_{i}\;{\rm d}t+\dfrac{1}{4\tau_{\mathrm{s}}}\sum_{j=1}^{N_{\mathrm{b}}}\left(\bm{D}_{i+1,j}-\bm{D}_{i,j}\right)\cdot\bm{F}_{j}^{\rm E}\;{\rm d}t+\sqrt{\dfrac{Q_{\mathrm{eq}}^{2}}{6\tau_{\mathrm{s}}}}\sum_{j=1}^{i+1}\left(\bm{B}_{i+1,j}-\bm{B}_{i,j}\right)\cdot{\rm d}\bm{W}_{j}(t).$ (2)
@@
The links between neighboring beads are phantom FENE springs with spring constant $H$. If $\zeta$ is the Stokes drag coefficient of the beads, then each link is characterized by an elastic time scale $\tau_{\mathrm{s}}=\zeta/4H$. The net spring force exerted on bead $i$ is $\bm{F}_{i}^{\rm E}=f_{i}\bm{Q}_{i}-f_{i-1}\bm{Q}_{i-1}$, where the FENE interaction coefficients $f_{i}=\left({1-|\bm{Q}_{i}|^{2}/Q_{\mathrm{m}}^{2}}\right)^{-1}$ ensure that the extension of each link does not exceed its maximum length $Q_{\mathrm{m}}$. The contour length or maximum length of the chain is then given by $R_{\mathrm{m}}=(N_{\mathrm{b}}-1)Q_{\mathrm{m}}$. [Note that in Eq. (2), one must set $\bm{Q}_{0}=\bm{Q}_{N_{\mathrm{b}}}=0$ in the equations for ${\rm d}\bm{Q}_{1}$} and ${\rm d}{\bm{Q}}_{N_{\mathrm{b}}-1}$.]
The Brownian forces that act on the beads are represented by independent increments of vectorial Wiener processes ${\rm d}\bm{W}_{i}(t)$. The competition between Brownian and elastic forces sets the equilibrium length scale $Q_{\mathrm{eq}}=\sqrt{3k_{B}T/H}$, where $k_{B}$ is the Boltzmann constant and $T$ is the temperature. $Q_{\mathrm{eq}}$ is the root-mean-square (r.m.s.) extension of the link at equilibrium in a still fluid, in the Hookean limit of $Q_{\mathrm{eq}}\ll Q_{\mathrm{m}}$. The corresponding r.m.s. end-to-end extension of the chain at equilibrium is $R_{\mathrm{eq}}=Q_{\mathrm{eq}}\sqrt{N_{\mathrm{b}}-1}$ [53].
@@
Equations (1) to (7) constitute the bead-spring chain model with hydrodynamic interactions. For a given number of beads, $N_{\mathrm{b}}$, the chain is parameterized by $Q_{\mathrm{eq}}$, $Q_{\mathrm{m}}$, $\tau_{\mathrm{s}}$, and the HI parameter $h^{\star}$ (setting $h^{\star}=0$ yields the free-draining Rouse model [53, 54]).
II.2 Dependence of chain parameters on the number of beads
In this work, we consider a polymer with a fixed contour length or maximum extension $R_{\mathrm{m}}$ and model it as a bead-spring chain, so that the number of beads represents the level of coarse-graining. To understand how the effect of HI depends on coarse-graining, we need a way to compare the results of chains with different numbers of beads. In the absence of HI, the mapping of Jin and Collins [60] has been found to work very well in homogeneous isotropic turbulent flows. Specifically, the PDF of $R/R_{\mathrm{eq}}$ of two chains with different values of $N_{\mathrm{b}}$ are found to be in agreement, provided the parameters are determined in the following manner. First, we fix the equilibrium length of the links $Q_{\mathrm{eq}}$. This choice also fixes the equilibrium length of the dumbbell version of the model, $R_{\mathrm{eq}}^{D}=Q_{\mathrm{eq}}$. Next, we choose values for the maximum extension and the elastic time scale for the dumbbell model: $R_{\mathrm{m}}^{D}$ and $\tau^{D}$. We then calculate the remaining link parameters, $Q_{\mathrm{m}}$ and $\tau_{\mathrm{s}}$, for a chain with $N_{\mathrm{b}}$ beads according to
$\frac{Q_{\mathrm{m}}}{Q_{\mathrm{eq}}}=\frac{1}{\sqrt{N_{\mathrm{b}}-1}}\,\frac{R_{\mathrm{m}}^{D}}{R_{\mathrm{eq}}^{D}},\quad\tau_{\mathrm{s}}=\frac{\quad 6\tau^{D}}{(N_{\mathrm{b}}+1)N_{\mathrm{b}}}.$ (8)
This mapping is consistent with the view that the chain is a coarse-grained representation of a polymer with a fixed number of Kuhn steps $N_{\textsc{k}}$. Recalling that ${N_{\textsc{k}}}^{1/2}$ is proportional to the ratio of the contour length of the chain $R_{\mathrm{m}}$ to its equilibrium r.m.s. extension $R_{\mathrm{eq}}$, we see that maintaining $N_{\textsc{k}}$ constant as we coarse-grain requires $R_{\mathrm{m}}/R_{\mathrm{eq}}$ to be constant. Now, for a fixed $Q_{\mathrm{eq}}$, we have $R_{\mathrm{eq}}=\sqrt{N_{\mathrm{b}}-1}Q_{\mathrm{eq}}$ (based on the random-walk theory for a Rouse chain [61, 53]). So, the mapping (8) yields $R_{\mathrm{m}}/R_{\mathrm{eq}}=(N_{\mathrm{b}}-1)Q_{\mathrm{m}}/\sqrt{N_{\mathrm{b}}-1}Q_{\mathrm{eq}}=R_{\mathrm{m}}^{D}/Q_{\mathrm{eq}}$, which is independent of $N_{\mathrm{b}}$ as we have fixed the values of both $R_{\mathrm{m}}^{D}$ and $Q_{\mathrm{eq}}$. Thus, the Jin-Collins mapping sets the link parameters so that the chain’s contour length $R_{\mathrm{m}}$ increases as $\sqrt{N_{\mathrm{b}}-1}$, thereby ensuring that $R_{\mathrm{m}}/R_{\mathrm{eq}}$ is independent of $N_{\mathrm{b}}$. Since $R_{\mathrm{m}}/R_{\mathrm{eq}}=R_{\mathrm{m}}^{D}/Q_{\mathrm{eq}}$, we see that the initial choices for $R_{\mathrm{m}}^{D}$ and $Q_{\mathrm{eq}}$ sets the extensibility of the polymer for all levels of coarse graining, including that of the dumbbell (for which $R_{\mathrm{eq}}^{D}=Q_{\mathrm{eq}})$. In our simulations, we set $Q_{\mathrm{eq}}=1$ and $R_{\mathrm{m}}^{D}=109.54$, yielding an extensibility ratio of ${N_{\textsc{k}}}=R_{\mathrm{m}}^{2}/R_{\mathrm{eq}}^{2}=12000$. While being relevant to long DNA molecules [23], which have been used in drag-reduction experiments [62, 63], this value of the extensibility ratio ensures a sufficient range of extension for a power-law distribution to manifest.
The time-scale $\tau^{D}$ in (8) may be thought of as an approximation to the longest relaxation time of the entire chain. Its relation to the time-scale of the individual springs $\tau_{\mathrm{s}}$ is consistent with the behavior of the largest relaxation time of a Rouse chain (no HI), which for large $N_{\mathrm{b}}$ scales as $N_{\mathrm{b}}^{2}\tau_{\mathrm{s}}$. The precise form of the relation between $\tau^{D}$ and $\tau^{s}$ was obtained by Jin and Collins [60] on substituting the relation between $R_{\mathrm{eq}}$ and $Q_{\mathrm{eq}}$ into an expression for the elongational viscosity of highly-stretched polymers, given by Wiest and Tanner [64], and requiring the resulting viscosity values to be independent of $N_{\mathrm{b}}$.
In the absence of HI, the Jin-Collins mapping has been found to work very well for bead-spring chains in homogeneous isotropic turbulent flows: on using the mapping, the PDF of $R/R_{\mathrm{eq}}$ of a dumbbell matches that of a chain with 10 or 20 beads. This is true, not only for the stationary PDF [48], but also for the evolving PDF that describes the initial stretching of polymers when they are first introduced into a turbulent flow [42]. Here, we use the Jin-Collins mapping as a basis of comparison that is particularly well-suited to revealing the effects of HI, given that it accounts for changes due to varying $N_{\mathrm{b}}$ in the absence of HI.
The value of $h^{\star}$ is, in principle, tied to the value of $\tau_{\mathrm{s}}$ through the bead radius $a$ [see (6) and recall that $\tau_{\mathrm{s}}=\zeta/4H$ where $\zeta=6\pi\mu a$, where $\mu$ is the solvent viscosity]. In practice, however, this relationship is often ignored and $h^{\star}$ is treated as a free parameter that controls the strength of HI [65, 11]; typically, its value is tuned to reproduce experimental observations, e.g., of the longest relaxation time [65]. Here, we vary $h^{\star}$ freely in order to understand the influence of HI on the stretching dynamics of the chain; specifically, we use $h^{\star}=0$ (free-draining), $h^{\star}=0.2$ (an intermediate and typical value [8, 9]), and $h^{\star}=0.49$ (the largest physically meaningful value).
II.3 Brownian dynamics simulations